The length of a playground exceeds twice its width by 25 feet, and 650 feet of fencing are needed to enclose it. Find its dimensions.
Answer by checkley77(12844)

L=2W+25
2L+2W=650
2(2W+25)+2W=650
4W+50+2W=650
6W=650-50
6W=600
W=600/6
W=100 FT. FOR THE WIDTH.
L=2*100+25)
L=2*100+25)
L=200+25
L=225 FT. FOR THE LENGTH.
PROOF:
2*225+2*100=650
450+200=650
650=650